jesternmais:

Juanita needed 3 and 2/3 hours to take a standardized test, and Jordan needed 5 and 1/4 hours. How much more time did Jordan need than Juanita to take the test?

jesternmais:

new question please help me

lilcutie2:

@AZ

AZ:

this is a simple subtraction question but with fractions and now I remember you asked the fraction questions in the past xD

AZ:

so we want to know how much more time Jordan needs than Juanita so are we doing 5 1/4 - 3 2/3 or 3 2/3 - 5 1/4

jesternmais:

this is a new question my teacher jus gave me and I don't understand

AZ:

remember Jordan needs more time and we want to know how much more time she takes compared to Juanita

jesternmais:

yea

AZ:

so what are we subtracting from what

jesternmais:

3 2/3 - 5 1/4

AZ:

mhmm are you sure? We want to know how much more Jordan did than Juanita so think about it like this I have 5 apples and you have 3 apples we want to know how many more apples I have compared to YOU so would we do 3 - 5 or would we do 5 - 3

jesternmais:

ok

jesternmais:

so it would be 5 - 3 right?

AZ:

yes!

AZ:

so look back at your question are we going to do \( 5\dfrac{1}{4} - 3\dfrac{2}{3}\) or \( 3 \dfrac{2}{3} - 5 \dfrac{1}{4}\)

jesternmais:

5 1/4−3 2/3

AZ:

good so now we have to make these fractions into improper fractions I helped you do that last week, do you still remember how?

jesternmais:

yea

jesternmais:

so you would add the whole numbers together which is 8

jesternmais:

no subtract

jesternmais:

which is 3

AZ:

no no, we can't subtract yet

AZ:

we have to make it an improper fraction first

jesternmais:

oh

jesternmais:

yea I forgot

AZ:

so remember 

AZ:

so that was an example but what is 5 1/4 as an improper fraction as one number all over 4

jesternmais:

5 1/4= 17/4

AZ:

are you sure mhmm what is 5 * 4 and then add 1

jesternmais:

2

jesternmais:

?

AZ:

5 times 4

jesternmais:

20

AZ:

and then add 1

jesternmais:

21

AZ:

so 5 1/4 is the same thing as 21/4 now what about 3 2/3 can you do the same thing?

jesternmais:

3 2/3 = 11/3

AZ:

good! so now we have 21/4 - 11/3 we need both fractions to have the same denominator so we can subtract them do you know how to make them have the same denominator?

jesternmais:

21/4 - 11/3 = 32/4

AZ:

no we are SUBRACTING not adding also we cannot subtract until we have the SAME denominator

jesternmais:

oh yea

jesternmais:

=10 /4

AZ:

like I said We need to have the SAME DENOMINATOR

AZ:

so let's make the denominator for both 12 so the first fraction is 21/4 we need to multiply 3 on both the numerator and denominator so we can get 12 as the denominator what is 21 * 3 = what is 4 * 3

jesternmais:

21 x 3=63

jesternmais:

4 x 3= 12

jesternmais:

now what do I do

AZ:

so 21/4 = 63/12 now what about 11/3 let's make the denominator 12 so we have to multiply top and bottom by 4 what is 11 * 4 and what is 3 * 4

jesternmais:

11 x 4 = 48

jesternmais:

3 x 4 = 12

jesternmais:

so 48/12

jesternmais:

so 48/12

AZ:

11 times 4 is not 48 please recheck that part

jesternmais:

44

AZ:

so we have 44/12 now altogether we have 63/12 - 48/12 can you subtract it now?

jesternmais:

yes

AZ:

the denominator is the same which is 12 so just subtract the top part (the numerator)

jesternmais:

ok

jesternmais:

ok

jesternmais:

63/12 - 48/12= 15/4

AZ:

the 15 is correct but the DENOMINATOR has to stay the SAME we had 12 and 12 so we can SUBTRACT and so your final answer has to be over 12

jesternmais:

yea 15/12*

AZ:

there you go! that's your answer

jesternmais:

thank you az

jesternmais:

thank you az

AZ:

you're most welcome!!